OpenStudy (cameronlovesmorgan):

In three more years, Angelo’s grandfather will be six times as old as Angelo was last year. When Angelo’s present age is added to his grandfather’s present age, the total is 68. How old is each person now?

OpenStudy (sachintha):

Form up two equations. Let age of Angelo be A and age of his grandfather be G. Then, \(\sf G+3 = 6(A-1)\) \(\sf A + G = 68\) Solving these equations you get the age of Angelo as 11 and age of his grandfather as 57.